# XAI-CF – Examining the Role of Explainable Artificial Intelligence in Cyber Forensics

### v2

@@ 4.2.1 Summary
In the growing field of XAI, the last decade has seen a significant surge in efforts aimed at unraveling the complexities of machine learning models. Researchers have delved into diverse domains, ranging from healthcare and finance to cybersecurity, employing Bayesian frameworks, optimization-based algorithms, and model-agnostic techniques. These approaches seek to strike a delicate balance between achieving high model accuracy and ensuring interpretability. By introducing methods such as LIME, SHAP, and Anchors, the research community is actively working towards demystifying the decision-making processes of complex models. These strides not only promise to enhance the overall performance of models but also address ethical concerns regarding transparency in decision systems, fostering a greater understanding and trust in the capabilities of advanced AI systems. These efforts are particularly evident in applications like healthcare monitoring, stock price prediction, and cybersecurity, where the integration of XAI elements such as LIME and SHAP offers transparent insights into model decisions. As XAI techniques continue to evolve, they play a pivotal role not only in optimizing model hyperparameters but also in elucidating feature importance and providing critical insights for decision-making in diverse and complex real-world scenarios. Ultimately, the advancements in XAI contribute to a more accountable and trustworthy integration of machine learning models into practical applications.
@@
[…]
@@
Technique
Articles
Types
I / P
G / L
S / A
LIME
(mishra2017local, ), (feichtner2020understanding, ), (szczepanski2021new, ), (gite2021explainable, )
P
L
A
SHAP
(gramegna2020buy, ), (parsa2020toward, ), (debjit2022improved, ), (nordin2023explainable, )
P
G
A
Grad-CAM
(joo2019visualization, ), (sutton2022artificial, ), (naeem2022explainable, )
P
L
A
Rule lists
(wang2015falling, ), (su2015interpretable, ), (letham2015interpretable, )
P
G
A
Anchors
(ribeiro2018anchors, ), (szczepanski2021new, )
P
L
A
Grammar evolution
(cavaliere2020parkinson, )
P
G
A
LORE
(guidotti2018local, )
P
L
A
Saliency map
(fong2017interpretable, )
P
L
A
Decision Trees
(hara2016making, )
P
G
S, A
Cluster-TREPAN
(de2020explainable, )
P
L
A
@@ 4.3 Cyber Forensics (CF)
A formal definition of CF is given in section 1.2. This section explores in more detail of why, what, and how of CF. Every contact by a perpetrator leaves behind traces (chisum2000evidence, ). To make a case against the perpetrator these traces or pieces of evidence need to be found, collected, secured, studied, and analyzed. Cyber Forensics (alam2022cybersecurity, ; sammons2012basics, ; garfinkel2010digital, ) uses scientific methods and expertise to gather and analyze pieces of evidence found in cyber devices that can be used in criminal or other investigations in a court of law. This evidence can be used for different purposes, such as electronic discovery, intelligence, and administrative. For example, the data collected from digital devices can provide actionable intelligence. This intelligence can help accomplish different types of missions, such as securing national interests, decreasing or eliminating crimes like kidnapping and child exploitation, etc. Electronic discovery is the process of searching, finding, and securing any electronic data later to be used in a civil or criminal forensic case.
The main purpose of a CF process is to better understand and make a conclusion about an event of interest by finding and analyzing the details and facts related to that event. In general this process comprises of four phases (kent2006guide, ) as shown in Figure 6 and explained below:
@@
    Collection – Evidence collection is the foremost piece of work during CF. While collecting the data there is a strong need of also preserve and secure that data. The evidence collected must be protected from accidental or intentional compromise. Cyber devices must be made inaccessible as soon as there is a guarantee that no volatile data will be lost. Collection is also performed promptly because of the dynamic nature of the data. The data may get lost in battery-powered devices and network connections if not collected promptly. Actions during the collection phase must be well documented. A forensic clone or image is made of the suspected cyber device and the exam is conducted on the cloned device. The chain of custody must be maintained and documented. A proper Chain of Custody is to be maintained to ensure that the evidence collected has not been altered or changed during the forensic process.
@@
    Examination – This is the process of forensically scrutinizing the data (small or large) using automated and manual methods/techniques. During this process data of particular interest is extracted and analyzed, meanwhile, preserving the integrity of the data. Moreover, the data files are defined that contain the information, which may be compressed, encrypted, and/or protected by access controls, which makes a meaningful extraction difficult. With the availability of AI techniques and tools sifting through and extracting useful data has become much easier. It also reduces the amount of data that needs to be analyzed. To trust such AI processes/techniques they need to be made explainable, and this makes the application of XAI in CF more compelling.
@@
    Analysis – In this phase results of the examination are analyzed, using legally justifiable methods and techniques. The purpose is to extract useful information from the data that can be used and presented in a court of law. Different techniques are used such as correlation of data, abnormality in data, such as slow network traffic, etc., and sifting through the data using automated (AI, ML, data mining, etc.) and manual techniques to predict an attack. Here the use of AI becomes extremely beneficial. Not only does it automate (for now partially) this process but also helps in extracting meaningful information from the data. But to accept such a piece of information in a court of law or public, there is a strong need to make the process of extracting such information explainable, and this makes the application of XAI in CF more indispensable.
@@
    Reporting – This is the last phase but also the most important phase of a forensics process. It makes the results of the analysis phase more comprehensive and understandable by both a technical and a non-technical person. It generally consists of the actions used during the forensic process, how and what tools/methods were used, and what other actions need to be performed, and may also provide recommendations and guidelines to improve the forensic process. It is challenging to provide an exact explanation but should provide as much detail as possible to the readers to help them make a responsible and accountable conclusion and decision about an event. Here XAI can play a major role in providing human understandable explanations that can be used by the forensic analyst to produce such a report.
CF is an important tool for solving crimes committed against people where the evidence resides on a cyber device (garfinkel2010digital, ). Initially, forensic techniques were developed primarily for recovering data. As the techniques developed network and memory forensics became possible and CF grew widespread and reliable. Now, with the rise of complex cyber devices such as smartphones, IoT, etc., the proliferation of operating systems and file formats, pervasive encryption, use of the cloud for remote processing and storage, and legal challenges, CF is facing many new challenges. For example, there are dozens of systems running on smartphones each with more than millions of downloadable applications. File carving is one of the techniques to extract and classify raw data from a cyber device, e.g., a smartphone, into files. Sifting through such a large amount of raw data for classification to make sense of this data requires new techniques, such as from the field of AI (alam2023sift, ). To apply these techniques successfully in CF we need to justify and explain the results to a human (hall2022explainable, ). Here the question arises does XAI provide such explainable results? This is the main focus of the study carried out in this paper.
@@ 4.3.1 Classifications
The basic CF techniques are the same but others are only applied in a certain environment, such as mobile, IoT, etc. Based on these working environments during the forensic process, in this section, we discuss some of the main classifications of CF. Every cyber device is kind of a computer device, i.e., it is capable of processing and storing electronic data. Therefore, we do not consider computer forensics as a separate class in CF. In this paper, we name this class as General and we do not explain this class exclusively here.
@@
    Mobile Forensics – Nowadays we cannot imagine our life without a smartphone. These devices are mini-computers capable of almost the same functionality as a desktop computer. The number of mobile devices is growing exponentially. Their hardware and operating systems (OSs) are also evolving rapidly. Evidence can be located not just on the device or memory but also on the cellular network. Any component (base station, mobile switching center, home, visitor location register, etc.) in the network can potentially provide information that can be used in a forensic investigation. OS is responsible for creating and storing evidence and has a significant impact on device forensics. There are several mobile OSs but the popular ones are Android, Apple, and Windows CE. Some of the potential items for collecting evidence in mobile devices are pictures and videos, deleted text messages, browser history, location information GPS, contacts, chat sessions, etc. The proliferation of mobile devices each with a different OS makes forensics of these devices challenging. Also, their migrating nature makes it difficult to find exactly the location of the device.
@@
    IoT Forensics – IoT devices are all over the Internet, such as in smart cities, healthcare, smart offices/homes, vehicle automation, etc. The IoT market will continue to experience exponential growth. Like mobile forensics IoT forensics is also facing a lot of challenges (stoyanova2020survey, ). A comprehensive list (nist2014nist, ) is provided by the National Institute of Standards and Technology (NIST). We discuss here some of the important challenges. Due to the highly dynamic nature of the communication, it is almost impossible to know the number of devices to be confiscated and the scope of the damage. It is quite possible that important data, such as registry entries or files, could be completely erased before it is collected. Like mobile devices, due to their moving/migrating nature locating the evidence becomes a challenge. The proliferation of these devices makes forensics of these devices difficult. In IoT, the hardware and software used are heterogeneous. Maintaining a proper chain of custody in IoT is difficult. The reason is the presence of evidence data on multiple remote servers as in the cloud. Some IoT devices have a very limited memory and can easily be overwritten resulting in the possibility of missing the evidence. One of the advantages of IoT forensics is that the evidence if mirrored to multiple locations as in the cloud is harder to destroy by the criminals.
@@
    Network Forensics – Application of scientific methods to investigate network breaches and vulnerabilities. These methods rely on discovering, capturing, identifying, and analyzing network traffic. Current network forensic mechanisms involve logging, packet marking, and heuristics based on various events of the network. Some of the major challenges of network forensics (khan2016network, ; qureshi2021analysis, ) are. Due to the high network speed, it is difficult to capture packets and also requires a large amount of space to store packets for later analysis. Network intruders use several techniques, such as spoofing, to hide their IP addresses. As such, identifying the original IP address of the attacker becomes very difficult for forensic investigators. Like IoT forensics, due to the virtualized and distributive nature of networks, it becomes difficult to identify the location of the network devices to extract data. Due to the frequent mobility of data the trust and integrity become low. That means the network may not have consistent, complete, and accurate data which becomes a major concern for forensic investigators. Cloud is a wide area network and is used to host users and resources that can communicate using cloud-based technologies. Cloud forensics is part of network forensics and faces similar challenges and we do not consider this as a separate class in CF.
@@
    Database Forensics – Any practical/working digital application contains a database of some kind. As an application becomes more complicated a database is used to store important and sensitive information. Database forensics deals with the detailed analysis of a database including its log files, metadata, and data files (chopade2019ten, ). Database forensics is an upcoming class of CF and not much work is done in this area. This is due to the inherent complexity and multidimensional nature of databases. Some of the major challenges of database forensics (al2014towards, ) are. One of the basic requirements to perform successful database forensics is the understanding of the internal architecture of the database. A variety of database infrastructures (Oracle, MySQL, SQLite, DB2, etc.) and platforms (Windows, Android, UNIX, Solaris, etc.) make this task very difficult. Multidimensional databases can be viewed from different perspectives/features. This simplifies data presentation, navigation, and maintenance, but also makes it difficult to forensically analyze the tampering of such data (fasan2012reconstruction, ). As in other classes, scanning and making sense of large data composed of data files, logs, and metadata pose a challenge to performing successful forensics. Moreover, determining the location of tampering and suitable data in this large set of data, which may contain manual vulnerabilities and weaknesses created by inexperienced database administrators/users, poses resource and time constraints.
@@ 4.4 Last Ten Years of CF
To make readers familiar with the study carried out in the paper, we present a critical and short review of the last ten years of work carried out in CF. We exclude papers that explore the application of XAI in CF because we discuss in detail such applications in section 6. A detailed survey on this topic is out of the scope of this paper. For a detailed survey, interested readers are referred to (casino2022research, ).
Neasbitt et al. (neasbitt2014clickminer, ) propose ClickMiner, a system designed to automatically reconstruct user-browser interactions from full packet network traces. Through a user study, the research demonstrates ClickMiner’s effectiveness in reconstructing interactions and highlights its potential for aiding forensic analysis of web traffic traces, particularly in investigating user-browser interactions before security incidents.Silde et al. (silde2014digital, ) develop and a methodology for profiling cyberstalkers by adapting behavior evidence analysis. This helps forensics investigators in tracing the activities of the cyberstalker and also reduces the time for digital investigation. Neasbitt et al. (neasbitt2015webcapsule, ) introduce a WebCapsule that presents a novel record and replay forensic engine for browsers, providing an always-on, lightweight solution for reconstructing web security incidents, especially phishing and social engineering attacks. It records non-deterministic inputs to the core web rendering engine, enabling detailed forensic analysis of unexpected incidents and ephemeral web content. Karpisek et al. (karpisek2015whatsapp, ) perform an analysis on WhatsApp’s calling feature, decrypting network traffic to extract key artifacts like phone numbers, server IPs, and call duration, etc. Their work also provides a visualization tool for providing essential insights for forensic examinations of network traffic related to potential suspects.
@@
Nalawade et al. (nalawade2016forensic, ) explore web browser forensics, focusing on collecting evidence from browsers like cache, history, cookies, and download lists. The study compares major tools for web browser analysis, emphasizing their benefits and limitations. Anglano et al. (anglano2016forensic, ) conduct a forensic analysis of ChatSecure, an Android Instant Messaging app, revealing vulnerabilities in its encrypted databases. Employing CF techniques, the study extracts the encryption passphrase from volatile memory, allowing decryption without user disclosure. Zhang et al. (zhang2017blockchain, ) present an approach, termed process provenance, to bolster the trustworthiness of the chain of custody in cloud forensics. By integrating blockchain and cryptography group signatures, the proposed method enhances proof of existence and privacy preservation for process records, addressing challenges arising from the distributed nature of cloud computing. Koroniotis et al. (koroniotis2019towards, ) introduce the Bot-IoT dataset, encompassing both normal and attack traffic, aiming to enhance the credibility of network intrusion detection and forensic systems. Statistical and machine learning methods are employed to evaluate the dataset’s reliability, showcasing its potential for botnet identification in IoT-specific networks and providing a foundation for future network forensic models. Nyaletey et al. (nyaletey2019blockipfs, ) introduce BlockIPFS, aiming to address traceability and security concerns in the Interplanetary File System (IPFS) by integrating blockchain technology. BlockIPFS, combines the efficiency of IPFS for file sharing with the traceability and security features of blockchain, offering a comprehensive solution for improved trustworthiness, authorship protection, and auditability in distributed file systems.
@@
Dalezios et al. (dalezios2020digital, ) focus on enhancing forensic capabilities in cloud platforms, addressing challenges tied to criminal activities. The Cloud Auditing Data Federation standard event model is employed to represent cloud-related events, promoting a clear and robust logging format for forensics investigations in cloud computing. Rahman et al. (rahman2020new, ) propose a web forensic framework that addresses the growing concern of automated bot crimes, providing a systematic approach. By leveraging real institutional web application logs and experimental bot attack scenarios, the authors develop unique access loggers in JavaScript and PHP to extract crucial patterns for forensic analysis. Noura et al. (noura2020distlog, ) propose a technique to countermeasure the anti-forensics methods that target the logs. The proposed technique first aggregates, compresses, and encrypts the logs. These logs are then fragmented, authenticated, and distributed over a number of storage nodes and secure the log files. Luo et al. (luo2020automatic, ) propose a framework for automatically inspecting and identifying Android applications that are inappropriate for kids under 12 years of age. Forensic information/content is extracted by parsing the HTTP packets. The contents are then classified as appropriate or inappropriate and a normal or abnormal report is generated, respectively. Hina et al. (hina2021sefaced, ) present SeFACED, an approach for email content classification. Outperforming traditional machine learning and deep learning models, SeFACED achieves robust accuracy and precision, contributing significantly to forensic investigations and enhancing the identification of malicious activities in email content analysis.
@@
Shahbazi et al. (shahbazi2022nlp, ) integrate natural language processing (NLP) techniques with a blockchain framework to enhance the security and performance of online social media investigations. By applying the Random Forest algorithm and blockchain technology, the study addresses system security and traces process changes, offering the potential for advancements in investigating cybercriminal activities and fraud. Kim et al. (kim2022methods, ) present a method for recovering deleted data from a Realm database. This database is mostly used and optimized for mobile devices. The method is based on database structure and type. Such a method can also be used against anti-forensics techniques, such as deleting evidential data from the database, used by adversaries. Gupta et al. (gupta2023digital, ) explore the digital artifacts left by the Discord social media application on the Google Chrome web browser, focusing on Windows 10. The study employs various techniques, including the examination of local storage, log files, and the Google Chrome cache, offering insights for digital forensic analysts and researchers in the domain of social media forensics. Nissan et al. (nissan2023database, ) present a framework for predicting database query activity from a memory snapshot. The authors use a Support Vector Machine for such prediction. This memory forensics allows forensic investigators to detect malicious activities even in the presence of incomplete logs. Dragonas et al. (dragonas2023iot, ) propose a forensic model to identify interconnectivity between IoT. IoT devices are connected to other IoT devices and identifying this connectivity of things is essential to conduct a thorough investigation. A tool is also presented for extracting and visualizing this connectivity. Domingues et al. (domingues2023post, ) present a study for identifying and analyzing CF artifacts available in a post-mortem examination of a popular Zepp Life (monitors health and other activities) application for Android. The CF artifacts from these devices have been used in various criminal investigations (domingues2023post, ).

### v3

@@ 4.2.1 Summary
Recent research shows that XAI is essential for improving transparency, trust, and reliability in fields such as healthcare, security, finance, the IoT, NLP, and multimodal deep learning. Many studies focus on developing explanation frameworks, like LIME, Anchors, and SHAP, to make complex models easier to understand and more useful. A key area of research adapts XAI to specific challenges. In healthcare, for instance, visual and feature-based explanations support medical diagnoses and COVID-19 screenings. In security, XAI helps detect threats and analyze malware. Tailored methods also apply to self-driving cars and stock prediction, demonstrating XAI’s ability to validate models and assist experts. Another important insight is that making models understandable can improve them. By visualizing key features or providing simple rules, studies often identify issues like model weaknesses and biases. This leads to stronger and more ethical models. User assessments show that the effectiveness of explanations relies on task complexity, model design, and user expertise. High-quality explanations need to be easy to understand, trustworthy, and practical. Overall, this review emphasizes progress in creating clear, reliable, and specialized AI systems, with hybrid methods and specific adaptations offering promising paths for future research.
@@
Elsevier13%PLOS5%IEEE17%Springer4%ACM5%ArXiv8%AAAI5%Nature17%PeerJ5%MDPI13%Hindawi4%ISMIR 4%
202513%202413%20234%202213%20219%202022%20194%20189%20179%2016 4%
@@
Technique Articles Types
I / P G / L S / A
SHAP [88], [90], [97], [98], [101], [99], [102], [103] P G A
LIME [83], [91], [93], [94], [101], [99], [104] P L A
Grad-CAM [87], [96], [95], [100] P L A
Saliency map [84], [100] P L A
Anchors [85], [93] P L A
Grammar evolution [89] P G A
LORE [86] P L A
Decision Trees [81] P G S, A
Cluster-TREPAN [92] P L A
@@ 4.3 Cyber Forensics (CF)
A formal definition of CF is given in section 1.2. This section explores in more detail of why, what, and how of CF. Every contact by a perpetrator leaves behind traces [105]. To make a case against the perpetrator, these traces or pieces of evidence need to be found, collected, secured, studied, and analyzed. Cyber Forensics [106, 107, 3] uses scientific methods and expertise to gather and analyze pieces of evidence found in cyber devices that can be used in criminal or other investigations in a court of law. This evidence can be used for different purposes, such as electronic discovery, intelligence, and administrative. For example, the data collected from digital devices can provide actionable intelligence. This intelligence can help accomplish different types of missions, such as securing national interests, decreasing or eliminating crimes like kidnapping and child exploitation, etc. Electronic discovery is the process of searching, finding, and securing any electronic data to be used in a civil or criminal forensic case.
The main purpose of a CF process is to better understand and draw a conclusion about an event of interest by finding and analyzing the details and facts related to that event. In general this process comprises of four phases [108] as shown in Figure 7 and explained below:
@@
    Collection – Evidence collection is the foremost piece of work during CF. While collecting the data, there is a strong need of also preserve and secure that data. The evidence collected must be protected from accidental or intentional compromise. Cyber devices must be made inaccessible as soon as there is a guarantee that no volatile data will be lost. Collection is also performed promptly because of the dynamic nature of the data. The data may get lost in battery-powered devices and network connections if not collected promptly. Actions during the collection phase must be well documented. A forensic clone or image is made of the suspected cyber device, and the exam is conducted on the cloned device. The chain of custody must be maintained and documented. A proper Chain of Custody is to be maintained to ensure that the evidence collected has not been altered or changed during the forensic process.
@@
    Examination – This is the process of forensically scrutinizing the data (small or large) using automated and manual methods/techniques. During this process, data of particular interest is extracted and analyzed, meanwhile preserving the integrity of the data. Moreover, the data files are defined that contain the information, which may be compressed, encrypted, and/or protected by access controls, which makes a meaningful extraction difficult. With the availability of AI techniques and tools, sifting through and extracting useful data has become much easier. It also reduces the amount of data that needs to be analyzed. To trust such AI processes/techniques, they need to be made explainable, and this makes the application of XAI in CF more compelling.
@@
    Analysis – In this phase results of the examination are analyzed, using legally justifiable methods and techniques. The purpose is to extract useful information from the data that can be used and presented in a court of law. Different techniques are used, such as correlation of data, abnormality in data, such as slow network traffic, etc., and sifting through the data using automated (AI, ML, data mining, etc.) and manual techniques to predict an attack. Here, the use of AI becomes extremely beneficial. Not only does it automate (for now, partially) this process, but it also helps in extracting meaningful information from the data. But to accept such a piece of information in a court of law or the public, there is a strong need to make the process of extracting such information explainable, and this makes the application of XAI in CF more indispensable.
@@
    Reporting – This is the last phase but also the most important phase of a forensics process. It makes the results of the analysis phase more comprehensive and understandable by both a technical and a non-technical person. It generally consists of the actions used during the forensic process, how and what tools/methods were used, and what other actions need to be performed, and may also provide recommendations and guidelines to improve the forensic process. It is challenging to provide an exact explanation, but one should provide as much detail as possible to the readers to help them make a responsible and accountable conclusion and decision about an event. Here, XAI can play a major role in providing human-understandable explanations that can be used by the forensic analyst to produce such a report.
CF is an important tool for solving crimes committed against people where the evidence resides on a cyber device [3]. Initially, forensic techniques were developed primarily for recovering data. As the techniques developed, network and memory forensics became possible, and CF grew widespread and reliable. Now, with the rise of complex cyber devices such as smartphones, IoT, etc., the proliferation of operating systems and file formats, pervasive encryption, use of the cloud for remote processing and storage, and legal challenges, CF is facing many new challenges. For example, there are dozens of systems running on smartphones, each with more than a million downloadable applications. File carving is one of the techniques to extract and classify raw data from a cyber device, e.g., a smartphone, into files. Sifting through such a large amount of raw data for classification to make sense of this data requires new techniques, such as from the field of AI [70]. To apply these techniques successfully in CF we need to justify and explain the results to a human [16]. Here, the question arises: Does XAI provide such explainable results? This is the main focus of the study carried out in this paper.
@@ 4.3.1 Classifications
The basic CF techniques are the same, but others are only applied in a certain environment, such as mobile, IoT, etc. Based on these working environments during the forensic process, in this section, we discuss some of the main classifications of CF. Every cyber device is kind of a computer device, i.e., it is capable of processing and storing electronic data. Therefore, we do not consider computer forensics as a separate class in CF. In this paper, we name this class as General and we do not explain this class exclusively here.
@@
    Mobile Forensics – Nowadays, we cannot imagine our life without a smartphone. These devices are mini-computers capable of almost the same functionality as a desktop computer. The number of mobile devices is growing exponentially. Their hardware and operating systems (OSs) are also evolving rapidly. Evidence can be located not just on the device or memory but also on the cellular network. Any component (base station, mobile switching center, home, visitor location register, etc.) in the network can potentially provide information that can be used in a forensic investigation. OS is responsible for creating and storing evidence and has a significant impact on device forensics. There are several mobile OSs, but the popular ones are Android, iOS, and Windows CE. Some of the potential items for collecting evidence in mobile devices are pictures and videos, deleted text messages, browser history, location information GPS, contacts, chat sessions, etc. The proliferation of mobile devices, each with a different OS, makes the forensics of these devices challenging. Also, their migrating nature makes it difficult to find the location of the device.
@@
    IoT Forensics – IoT devices are all over the Internet, such as in smart cities, healthcare, smart offices/homes, vehicle automation, etc. The IoT market will continue to experience exponential growth. Like mobile forensics, IoT forensics is also facing a lot of challenges [109]. A comprehensive list [110] is provided by the National Institute of Standards and Technology (NIST). We discuss here some of the important challenges. Due to the highly dynamic nature of the communication, it is almost impossible to know the number of devices to be confiscated and the scope of the damage. It is quite possible that important data, such as registry entries or files, could be completely erased before it is collected. Like mobile devices, due to their moving/migrating nature, locating the evidence becomes a challenge. The proliferation of these devices makes the forensics of these devices difficult. In IoT, the hardware and software used are heterogeneous. Maintaining a proper chain of custody in IoT is difficult. The reason is the presence of evidence data on multiple remote servers as in the cloud. Some IoT devices have a very limited memory and can easily be overwritten, resulting in the possibility of missing the evidence. One of the advantages of IoT forensics is that the evidence, if mirrored to multiple locations, as in the cloud, is harder to destroy by the criminals.
@@
    Network Forensics – Application of scientific methods to investigate network breaches and vulnerabilities. These methods rely on discovering, capturing, identifying, and analyzing network traffic. Current network forensic mechanisms involve logging, packet marking, and heuristics based on various events of the network. Some of the major challenges of network forensics [111, 112] are. Due to the high network speed, it is difficult to capture packets and also requires a large amount of space to store packets for later analysis. Network intruders use several techniques, such as spoofing, to hide their IP addresses. As such, identifying the original IP address of the attacker becomes very difficult for forensic investigators. Like IoT forensics, due to the virtualized and distributed nature of networks, it becomes difficult to identify the location of the network devices to extract data. Due to the frequent mobility of data, the trust and integrity become low. That means the network may not have consistent, complete, and accurate data, which becomes a major concern for forensic investigators. Cloud is a wide area network and is used to host users and resources that can communicate using cloud-based technologies. Cloud forensics is part of network forensics and faces similar challenges and we do not consider this as a separate class in CF.
@@
    Database Forensics – Any practical/working digital application contains a database of some kind. As an application becomes more complicated, a database is used to store important and sensitive information. Database forensics deals with the detailed analysis of a database, including its log files, metadata, and data files [113]. Database forensics is an upcoming class of CF, and not much work has been done in this area. This is due to the inherent complexity and multidimensional nature of databases. Some of the major challenges of database forensics [114] are. One of the basic requirements to perform successful database forensics is the understanding of the internal architecture of the database. A variety of database infrastructures (Oracle, MySQL, SQLite, DB2, etc.) and platforms (Windows, Android, UNIX, Solaris, etc.) make this task very difficult. Multidimensional databases can be viewed from different perspectives/features. This simplifies data presentation, navigation, and maintenance, but also makes it difficult to forensically analyze the tampering of such data [115]. As in other classes, scanning and making sense of large data composed of data files, logs, and metadata pose a challenge to performing successful forensics. Moreover, determining the location of tampering and suitable data in this large set of data, which may contain manual vulnerabilities and weaknesses created by inexperienced database administrators/users, poses resource and time constraints.
@@ 4.4 Last Ten Years of CF
To make readers familiar with the study carried out in the paper, we present a critical and short review of the last ten years of work carried out in CF. We exclude papers that explore the application of XAI in CF because we discuss in detail such applications in section 6. A detailed survey on this topic is out of the scope of this paper. For a detailed survey, interested readers are referred to [116].
@@
Nalawade et al. [117] explore web browser forensics, focusing on collecting evidence from browsers like cache, history, cookies, and download lists. The study compares major tools for web browser analysis, emphasizing their benefits and limitations. Anglano et al. [118] conduct a forensic analysis of ChatSecure, an Android Instant Messaging app, revealing vulnerabilities in its encrypted databases. Employing CF techniques, the study extracts the encryption passphrase from volatile memory, allowing decryption without user disclosure. Zhang et al. [119] present an approach, termed process provenance, to bolster the trustworthiness of the chain of custody in cloud forensics. By integrating blockchain and cryptography group signatures, the proposed method enhances proof of existence and privacy preservation for process records, addressing challenges arising from the distributed nature of cloud computing. Koroniotis et al. [120] introduce the Bot-IoT dataset, encompassing both normal and attack traffic, aiming to enhance the credibility of network intrusion detection and forensic systems. Statistical and machine learning methods are employed to evaluate the dataset’s reliability, showcasing its potential for botnet identification in IoT-specific networks and providing a foundation for future network forensic models. Nyaletey et al. [121] introduce BlockIPFS, aiming to address traceability and security concerns in the Interplanetary File System (IPFS) by integrating blockchain technology. BlockIPFS combines the efficiency of IPFS for file sharing with the traceability and security features of blockchain, offering a comprehensive solution for improved trustworthiness, authorship protection, and auditability in distributed file systems.
@@
Dalezios et al. [122] focus on enhancing forensic capabilities in cloud platforms, addressing challenges tied to criminal activities. The Cloud Auditing Data Federation standard event model is employed to represent cloud-related events, promoting a clear and robust logging format for forensic investigations in cloud computing. Rahman et al. [123] propose a web forensic framework that addresses the growing concern of automated bot crimes, providing a systematic approach. By leveraging real institutional web application logs and experimental bot attack scenarios, the authors develop unique access loggers in JavaScript and PHP to extract crucial patterns for forensic analysis. Noura et al. [124] propose a technique to countermeasure the anti-forensics methods that target the logs. The proposed technique first aggregates, compresses, and encrypts the logs. These logs are then fragmented, authenticated, and distributed over several storage nodes and secure the log files. Luo et al. [125] propose a framework for automatically inspecting and identifying Android applications that are inappropriate for kids under 12 years of age. Forensic information/content is extracted by parsing the HTTP packets. The contents are then classified as appropriate or inappropriate and a normal or abnormal report is generated, respectively. Hina et al. [126] present SeFACED, an approach for email content classification. Outperforming traditional machine learning and deep learning models, SeFACED achieves robust accuracy and precision, contributing significantly to forensic investigations and enhancing the identification of malicious activities in email content analysis.
@@
Shahbazi et al. [127] integrate natural language processing (NLP) techniques with a blockchain framework to enhance the security and performance of online social media investigations. By applying the Random Forest algorithm and blockchain technology, the study addresses system security and traces process changes, offering the potential for advancements in investigating cybercriminal activities and fraud. Kim et al. [128] present a method for recovering deleted data from a Realm database. This database is mostly used and optimized for mobile devices. The method is based on the database structure and type. Such a method can also be used against anti-forensics techniques, such as deleting evidential data from the database, used by adversaries. Gupta et al. [129] explore the digital artifacts left by the Discord social media application on the Google Chrome web browser, focusing on Windows 10. The study employs various techniques, including the examination of local storage, log files, and the Google Chrome cache, offering insights for digital forensic analysts and researchers in the domain of social media forensics. Nissan et al. [130] present a framework for predicting database query activity from a memory snapshot. The authors use a Support Vector Machine for such a prediction. This memory forensics allows forensic investigators to detect malicious activities even in the presence of incomplete logs. Dragonas et al. [131] propose a forensic model to identify interconnectivity between IoT. IoT devices are connected to other IoT devices, and identifying this connectivity of things is essential to conducting a thorough investigation. A tool is also presented for extracting and visualizing this connectivity. Domingues et al. [132] present a study for identifying and analyzing CF artifacts available in a post-mortem examination of a popular Zepp Life (monitors health and other activities) application for Android. The CF artifacts from these devices have been used in various criminal investigations [132].
Mei et al. [133] introduce a network-forensics framework that applies deep learning to attribute Advanced Persistent Threats (APTs)—that is, to infer an attacker’s identity, group affiliation, or campaign connections—from network telemetry. They aim to advance beyond simple detection and toward richer attribution and contextual forensic intelligence by learning patterns from large-scale traffic data, flow characteristics, and temporal behaviors. Alqahtany et al. [134] present an integrated blockchain and deep-learning model designed to enhance forensic traceability and preserve evidence integrity in Mobile VPN environments. Blockchain is employed to provide tamper-resistant logging, chain-of-custody verification, and distributed provenance, while the deep-learning component evaluates Mobile-VPN traffic and device artifacts to flag anomalies, identify misuse, and reconstruct session activity. Sharma et al. [135] propose ForensicLLM, a locally deployable large language model tailored for digital-forensics workflows. In contrast to cloud-based LLMs, ForensicLLM focuses on privacy, offline usability, reproducibility, and specialized reasoning for tasks such as log examination, artifact interpretation, and incident triage. Kao et al. [136] describe an NLP-driven system aimed at detecting cryptocurrency seed phrases (mnemonics) across multilingual datasets. Because such mnemonics appear in diverse linguistic forms, conventional pattern-matching techniques are insufficient; the authors therefore apply language-modeling methods to automatically identify, categorize, and extract mnemonic phrases within text sources encountered during cybercrime investigations.